Vaughn Corporation manufactures a single product. The standard cost per unit of product is shown below.

Direct materials—1 pound plastic at $7.00 per pound $ 7.00
Direct labor—1.5 hours at $11.10 per hour 16.65
Variable manufacturing overhead 11.25
Fixed manufacturing overhead 6.75
Total standard cost per unit $41.65


The predetermined manufacturing overhead rate is $12.00 per direct labor hour ($18.00 ÷ 1.5). It was computed from a master manufacturing overhead budget based on normal production of 8,850 direct labor hours (5,900 units) for the month. The master budget showed total variable costs of $66,375 ($7.50 per hour) and total fixed overhead costs of $39,825 ($4.50 per hour). Actual costs for October in producing 4,700 units were as follows.

Direct materials (4,890 pounds) $ 35,208
Direct labor (6,900 hours) 78,660
Variable overhead 62,602
Fixed overhead 24,098
    Total manufacturing costs $200,568


The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored.

(a)

Compute all of the materials and labor variances.

Total materials variance $

Neither favorable nor unfavorableUnfavorableFavorable

Materials price variance $

Neither favorable nor unfavorableUnfavorableFavorable

Materials quantity variance $

Neither favorable nor unfavorableUnfavorableFavorable

Total labor variance $

Neither favorable nor unfavorableUnfavorableFavorable

Labor price variance $

UnfavorableNeither favorable nor unfavorableFavorable

Labor quantity variance $

Neither favorable nor unfavorableFavorableUnfavorable


(b)

Compute the total overhead variance.

Total overhead variance $

FavorableNeither favorable nor unfavorableUnfavorable
Answer #1

solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
4890 $7.20 $35,208.00 4890 $7.00 $34,230.00 4700 $7.00 $32,900.00
$978.00 Unfavorable $1,330.00 Unfavorable
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $978.00 Unfavorable
Direct material quantity variance $1,330.00 Unfavorable
Total direct material variance $2,308.00 Unfavorable
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
6900 $11.40 $78,660.00 6900 $11.10 $76,590.00 7050 $11.10 $78,255.00
$2,070.00 Unfavorable $1,665.00 Favorable
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $2,070.00 Unfavorable
Direct Labor Efficiency variance $1,665.00 Favorable
Total direct labor variance $405.00 Unfavorable

solution b:

Total overhead variance = Overhead applied - Actual overhead

= (4700*1.5*$12) - ($62,602 + $24,098)

= $84,600 - $86,700 = $2,100 Unfavorable